OpenStudy (anonymous):

A square poster has sides measuring 2 feet less than the sides of a square sign. If the difference between their areas is 28 square feet, find the lengths of the sides of the poster and the sign. Can someone just set me up with an equation to solve?

OpenStudy (anonymous):

Sure thing::  The drawing is more for me than for you: Area = (x-2)(x-2) and x^2 The x^2 is going to be bigger because each side is more so: x^2-(x-2)(x-2) =28

OpenStudy (anonymous):

thank u but wat would be the next step if you could guide

OpenStudy (anonymous):

Well, next you want to factor that (x-2)(x-2) to get x^2-4x+4. Put that into your equation and solve for x.

OpenStudy (anonymous):

Am sorry am just so confused

OpenStudy (anonymous):

Okay: From the top: Do you get my drawing? The part where the one side is x and the other is x-2? And then do you know how to find the area of each?

OpenStudy (anonymous):

I understand the first part its just that first of all am not food with word probleams and second I dont remember how to do are with x I only was thought with numbers so you must understand the major confusion and stuggle am having

OpenStudy (anonymous):

Do you get how I found my equation of:\[x ^{2}-(x-2)(x-2)=28?\]

OpenStudy (anonymous):

yes i do am just not sure if I am supost to -28 to get 0 on one side or if am just confusing my self

OpenStudy (anonymous):

Well, this problem actually works out nicely but first we need to factor that (x-2)(x-2). (x-2)(x-2)=x^2-4x+4 \\We can go over this? Yes/No? So x^2-x^2+4x-4=28 Simplify that.

OpenStudy (anonymous):

Am ok with the equation its just the figuring out part that is confusing me because of the x

OpenStudy (anonymous):

We are trying to discover what x is. x is a variable, so in free space it can be any thing. But we have an equation for it: x^2-x^2+4x-4=28 Solving that equation will give you x. The x^2's cancel out and you get 4x-4=28 What is x?

OpenStudy (anonymous):

x=8

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Yes. But don't forget, that is only have of the problem. You also need to know what the length of the side of the poster is.

OpenStudy (anonymous):

so do i plug in 8 in the equation to get the side of the poster

OpenStudy (anonymous):

What is your equation?

OpenStudy (anonymous):

thank u with ur help i got my answer